OpenStudy (anonymous):

A particle moves along a line so that at time t, its position is s= 4 sin 4t. a) when does the particle change direction?

OpenStudy (zzr0ck3r):

if we assume t = 0 is the start, then it will change when we get the max value of 4sin(4t) we know the max value of sin(x) = 1 so max 4sin(4t) = 4 solve this sin(4t) = 1 this happens when 4t = pi/2 so t= pi/8

OpenStudy (anonymous):

thanks a lot!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

what I showed will give you the first time it changes direction this is every time it changes direction \[4t=\frac{\pi}{2}+\pi*k\] where K is a whole number \[t=\frac{\pi}{8}+\frac{\pi}{4}k\]

OpenStudy (zzr0ck3r):

do you understand all the steps?

OpenStudy (anonymous):

wait.. i will try to do it myself..

OpenStudy (anonymous):

I understand.. but the correct answer are \[t=\frac{ \pi }{ 4 }+\pi K, t=\frac{ 3\pi }{ 4 }+\pi K \] for positive integers K

OpenStudy (zzr0ck3r):

hmm i guess it changes directions when 4sin(4t) = 0?

OpenStudy (zzr0ck3r):

that does not make sense to me. sorry I told you the wrong think...